- Shoda left the TMS damped in 6DOFs with LVDTs over the weekend. It successfully continued damping. (Spectrograms are to be updated)
AllControlledSpectra190520.PNG: The example of the spectra. The peak at around 1.3 Hz remains, which may be because I decreased the control gain for L and T in order to decrease the induced control noise.